# History of $N(1680)$

### v2

@@ 4 Conclusion
What would Mitya, Vitya, and Maxim make of the current situation in particle physics? On the one hand, they would be justifiably proud of the robust successes of an instanton vacuum approach, which did so much to place it on a firm footing. On the other hand, some of the problems that concerned them remain unresolved, notably the exotic $\Theta^{+}$ and its partners.
Actually, twenty years ago, in 2003, two experimental groups, LEPS and DIANA, announced the observation of a light and narrow exotic baryon with mass in the vicinity of the $1540~{}\mathrm{MeV}$, which was predicted by our friends. The history of this discovery and its theoretical interpretations can be found in the recent review [28].
@@ Acknowledgment
My archive kept a letter for the name $\Theta^{+}$’ that Mitya sent to the community [29]. The result is that the community accepted Mitya’s name.
@@
To: Takashi Nakano, nakano@rcnp.osaka-u.ac.jp,

### v3

@@ 4 Conclusion
What would Mitya, Vitya, and Maxim think of the current situation in particle physics? On the one hand, they would be justifiably proud of the robust successes of an instanton-vacuum approach, which did so much to place it on a firm footing. However, some of the problems that concern them remain unresolved, notably the exotic $\Theta^{+}$ and its partners.
In fact, twenty years ago, in 2003, two experimental groups, LEPS and DIANA, announced the observation of a light and narrow exotic baryon with mass in the vicinity of $1540~{}\mathrm{MeV}$, which was predicted by our friends. The history of this discovery and its theoretical interpretations can be found in the recent review [30].
Often memory does not keep details, and let me remind two events associated with Mitya, Vitya, and Maxim and $\Theta^{+}$
1) In 2009, I organized a Workshop on Narrow Nucleon Resonances: Predictions, Evidences, Perspectives, Edinburgh, Scotland, June, 2009.
https://gwdac.phys.gwu.edu/ igor/Edinburgh2009/index.htm . Unfortunately all three friends (Maxim was one of organizers) missed this event for some reasons, but many people attended and discussed physics.
2) Back in 2013, Nara, Japan, hosted the XVth International Conference on Hadron Spectroscopy (Hadron2013), Nara, Japan, in November 2013, which had a Memorial section for Mitya, https://pos.sissa.it/205/. Vitya was there. The contribution of Yakov Azimov and myself is in Ref. [31].
Finally, a photo taken in Maxim’s apartment in Bochum, Germany, back to 2006 (Fig. 8).
Figure 8: Sitting on sofa from left to right: Masha, Maxim’s daughter, me, Maxim, Takashi Nakano. Slava Kuznetsov is sitting just in front of us, he was a chef of our dinner.
@@ Acknowledgment
My archive kept a letter for the name $\Theta^{+}$’ that Mitya sent to the community [32]. The result is that the community accepted Mitya’s suggestion.
@@
To: Takashi Nakano,